OpenStudy (anonymous):

F(x)=x2-8x+2 vertex? Again, I used -b/2a but it wasn't right ! What am I doing wrong??

OpenStudy (anonymous):

ok in this case a = 1 and b = -8 so \[-\frac{b}{2a}=-\frac{-8}{2}=4\]

OpenStudy (anonymous):

plug in 4 for x get \[4^2-8\times 4 + 2=16-32+2=-16+2=-14\]

OpenStudy (anonymous):

vertex is (4,-14)

OpenStudy (anonymous):

now sure what you did wrong but my guess is that you may have put \[-\frac{b}{2a}=-4\] but there are two minus signs (one in the formula and one from the -8)

OpenStudy (anonymous):

Thank you again! My teacher is horrible at this, se explains nothing!

OpenStudy (anonymous):

hope this helps